A thin uniform rod (mass = 0.440 kg) swings about an axis that passes through one end of the rod and is perpendicular to the plane of the swing. The rod swings with a period of 1.65 s and an angular amplitude of 10.2°. What is the length of the rod? What is the maximum kinetic energy of the rod as it swings?
Answer #1

The period is given as

T = 2 \pi \sqrt{\frac{I}{m g h}}

I is the moment of inertia at one of the rod I = (1.3) mL^2 and h = L/2 so

T = 2 \pi \sqrt{\frac{\frac{1}{3}mL^2}{m g \frac{L}{2}}}

L = \frac{3gT^2}{8 \pi^2}

L = \frac{3 (9.8)(1.65)^2}{8 \pi^2} = 1.014m

b) the maximum kinetic energy is equal to the potential energy at the highest point (as per center of mass) as energy is conserved in the pendulum

U = m g h = m g (L/2) (1 - \cos \theta)

U = (0.44) (9.8) (1.014/2) (1 - \cos 10.2)

U = 0.03455 J